# Probabilistic metaplasticity for continual learning with memristors

### v2

@@ Abstract
Edge devices operating in dynamic environments critically need the ability to continually learn without catastrophic forgetting. The strict resource constraints in these devices pose a major challenge to achieve this, as continual learning entails memory and computational overhead. Crossbar architectures using memristor devices offer energy efficiency through compute-in-memory and hold promise to address this issue. However, memristors often exhibit low precision and high variability in conductance modulation, rendering them unsuitable for continual learning solutions that require precise modulation of weight magnitude for consolidation. Current approaches fall short to address this challenge directly and rely on auxiliary high-precision memory, leading to frequent memory access, high memory overhead, and energy dissipation. In this research, we propose probabilistic metaplasticity, which consolidates weights by modulating their update probability rather than magnitude. The proposed mechanism eliminates high-precision modification to weight magnitudes and, consequently, the need for auxiliary high-precision memory. We demonstrate the efficacy of the proposed mechanism by integrating probabilistic metaplasticity into a spiking network trained on an error threshold with low-precision memristor weights. Evaluations of two continual learning benchmarks show that probabilistic metaplasticity achieves state-of-the-art performance while consuming $\sim 67\%$ lower memory for additional parameters and up to two orders of magnitude lower energy during parameter updates compared to an auxiliary memory-based solution. The proposed model shows potential for energy-efficient continual learning with low-precision emerging devices.
@@ Introduction
Autonomous edge devices constitute a major application domain of artificial intelligence–ranging from indoor robots to autonomous drones/vehicles [1]. These devices often need on-device training for real-time adaptation to dynamic environments [2]. However, training neural networks requires energy-intensive data movement and memory accesses, which can strain the energy budget of these resource-constrained platforms [3]. The challenge exacerbates in real-world scenarios where data is non-stationary and does not maintain independent and identical distribution. Conventional artificial intelligence fails to learn effectively in this scenario and catastrophically forgets previously learned information [4]. The ability to learn and adapt to non-stationary data streams requires continual learning mechanisms, which entail additional memory and computational overhead, further straining the strict resource budget on edge devices [5]. The constraints in autonomous edge applications present a unique set of challenges, which lead to certain desirable criteria for continual learning solutions. First, the continual learning solution should be task-agnostic, i.e., it should be effective with no supervision or task-boundary awareness for autonomous applications. Second, edge devices have limited on-chip memory, which prohibits numerous iterative training over a large dataset. Hence, these devices should learn in an online scenario where data are encountered only once [1]. Finally, the memory and computational overhead of the continual learning mechanism should be minimized considering the strict energy constraints of edge devices [6].
@@
Although spiking models incorporated in memristor crossbar architectures are promising, they show non-ideal characteristics which can limit their suitability for continual learning mechanisms. The weight consolidation mechanism mentioned above is a prominent method of continual learning that relies on precise modulation of the weight magnitude. This is extremely challenging to achieve with memristors since these nanoscale devices show low precision ($\sim 3-6$ bits [20, 21]) and inherent variability during conductance modulation. Recent research has circumvented this issue by incorporating auxiliary high-precision memory (usually implemented with SRAM) alongside memristor crossbars [22]. Binary memristor devices have been used to accelerate forward pass in continual learning tasks while training with backpropagation is carried out in high-precision memory [23]. Memristors with multi-level conductance have also been utilized for continual learning with metaplasticity, where auxiliary high-precision memory accumulates weight gradients, and memristor weights are updated when the accumulated gradient is equivalent to the device conductance resolution [24]. These approaches have shown promising performance with low-precision memristor weights in continual learning benchmarks. However, these solutions rely on multi-epoch training and knowledge of task progression in order to either assign separate output neurons for different tasks or to modulate hyperparameters. These prerequisites are not suitable for autonomous applications. Moreover, high-precision memory with SRAMs lead to a large silicon footprint and idle standby leakage power (tens to hundreds of picoWatts per bit) [25]. Finally, while crossbar architectures with auxiliary high-precision memory retain energy efficiency during inference, excessive memory access during training can potentially void the benefits of the compute-in-memory architecture.
@@
In this research, we explore a novel weight consolidation mechanism suitable for low-precision memristor weights. Neuroscience literature has shown that binary metaplastic synapses characterized by probabilistic transitions help memory retention [26]. Inspired by this, we propose the probabilistic metaplasticity model, where a weight’s plasticity translates to its probability of update rather than its magnitude of update. The probability of a weight’s update evolves during training such that important weights are assigned lower update probability to prevent catastrophic forgetting. Since this model does not need precise modulation in weight magnitude, the need for auxiliary high-precision memory is eliminated. Previous research explored probabilistic metaplasticity in the context of binary synapses in models of memory [26] and binarized networks that only support training a single-layer network [27]. We generalize this concept to memristor weights with multi-level conductance incorporated in a multilayer spiking network trained based on an error threshold. We evaluate the proposed solution on two image detection continual learning benchmarks considering single and multi-memristor weights. Evaluations show that low-precision memristor weights with probabilistic metaplasticity can achieve performance equivalent to state-of-the-art networks with high-precision weights in the challenging online task-agnostic continual learning scenario. Compared to an equivalent solution with high-precision auxiliary memory, probabilistic metaplasticity reduces memory overhead for additional parameters by $\sim 67\%$ while requiring up to two orders of magnitude lower energy during the parameter update phase. The proposed model is promising for memory- and energy-efficient continual learning for autonomous edge applications.
@@
Figure 1: Probabilistic metaplasticity with error threshold-based training. a. Spiking network trained on error threshold where the network weights are realized with 1T1R memristor crossbar array. The dendritic compartments in the hidden and output layer neurons integrate the error. When the error reaches the threshold $U_{\text{th}}$, the memristor weights are updated to the next higher (negative error) or lower (positive error) conductance level with probability $p_{\text{update}}$. b. Mean and standard deviation of resistance levels versus programming compliance current for the 1T1R memristor device (shown in inset) adopted in this work [20]. c. Update probability of weights for different values of the metaplastic coefficient $m$ and the weights $w$. $m$ is directly proportional to the activity level of adjacent neurons. We see that weights with highly active adjacent neurons (high $m$) and weight magnitude (high $|w|$) lead to low update probability.
@@
Here, $m_{ij}$ is a metaplastic coefficient assigned to each weight. As the network learns, the metaplastic coefficient evolves based on the activity trace of the adjacent neurons as -
$m_{ij}(t+1)=\begin{cases}m_{ij}(t)+{\Delta}m,&\textrm{\leavevmode\nobreak\ if}\>X^{\text{tr}}_{j}\geq m^{\text{pre}}_{\text{th}}\&X^{\text{tr}}_{i}\geq m^{\text{post}}_{\text{th}}\\ m_{ij}(t)&\textrm{\leavevmode\nobreak\ otherwise}\end{cases}.$ (2)
We see that high activity in the adjacent neurons increases the metaplastic coefficient of a weight. Figure 1c shows the update probability for different values of weights and metaplastic coefficients. We see that weights with higher magnitude and metaplastic coefficient exhibit lower update probability. Throughout training, a weight’s magnitude and metaplastic coefficient adapt in response to the activity of the network. This changes the update probability/plasticity of a weight, leading to a metaplastic effect to consolidate weights.
@@
                                     if $p_{\text{update}}<\epsilon$ then
@@
            Update metaplastic coefficient: $m\leftarrow m+{\Delta}m$
@@
Algorithm 1 Probabilistic metaplasticity with error threshold-based eRBP for continual learning in a spiking network. $\mathcal{T}$ denotes a set of sequential tasks. $S^{\text{in}}$ and $S^{\text{out}}$ are the input and output spike train and $W$ denotes weights realized with memristors. $U$ is the error accumulated at the dendritic compartment of neurons. When |$U$| crosses the error threshold $U_{\text{th}}$, the update probability $p_{\text{update}}$ of eligible weights are calculated and compared with a random number to decide which weights should be updated. The memristor weights selected for update are programmed to the next higher or lower conductance level depending on the sign of the error. The function Program() refers to the operations required to update a memristor’s conductance. Details of the eRBP algorithm can be found in methods.
@@ Results
We first evaluate probabilistic metaplasticity on the split-MNIST benchmark, which consists of 5 sequential image classification tasks, each composed of two classes from the MNIST dataset [34]. We conduct network-level simulations in Python considering a spiking network with 200 hidden neurons and 2 output neurons. The simulation emulates in situ learning with memristor weights and captures the quantized nonuniform distribution and the associated variance of memristor conductance during weight updates. Figure 2 shows the evolution of the network performance in each of the benchmark tasks as the network trains on them sequentially. When trained with error threshold-based eRBP with no probabilistic metaplasticity, we see that the network learns each task it encounters with high accuracy (Figure 2a). However, after training on all tasks, it retains high accuracy only on the recently encountered tasks, and the earlier task accuracies drop to random guess level.
Figure 2: Probabilistic metaplasticity with error threshold-based training for memristor weights. Evolution of task accuracies as the network is trained sequentially on the split-MNIST tasks. The x-axis shows the latest task the network has learned. We evaluate the performance of task $n$ only after the network encounters it and denote the accuracies as 0 before that. a. With no probabilistic metaplasticity, the network learns the current task well, but forgets the initial tasks after sequentially learning multiple tasks. b. Probabilistic metaplasticity with low activity threshold leads to high rigidity in the network, so it remembers previous tasks but cannot learn the last task. c. High activity threshold can lead to loss in previous task accuracy while the network remains plastic to learn the new tasks. d. Optimized activity threshold balances plasticity and rigidity such that the network maintains high initial task accuracies while maintaining the ability to learn new tasks.
@@
We further evaluate probabilistic metaplasticity on the split-Fashion MNIST benchmark, which also consists of five sequential image classification tasks, each with two classes [35]. To observe the effect of weight resolutions, we consider both single- and multi-memristor weights[36]. Multi-memristor weights are configured by connecting $n_{\text{mem}}$ number of memristors in parallel. Figure 1a shows a 1T1R crossbar with multi-memristor weights consisting of three 1T1R devices. During training, only one memristor arbitrated by a global counter is updated to reduce training overhead. The parallel combination of devices increases the dynamic range of the weight conductance by a factor of $n_{\text{mem}}$ and increases the weight resolution. Table 1 lists the mean accuracy across benchmark tasks after sequential training for probabilistic metaplasticity and state-of-the-art weight consolidation-based continual learning models in networks with full-precision weights. We also list the memory required by each model for parameters in addition to weights (details in Supplementary note S1). The baseline results are for a spiking network with full-precision weights trained with eRBP. We note an improvement in the mean accuracy across tasks with increased weight resolution, with the best performance when $n_{\text{mem}}$ = 7, leading to a mean weight resolution $\sim 6$-bit. Results show that with probabilistic metaplasticity, a spiking network with noisy low-precision multi-memristor weights achieves performance equivalent to state-of-the-art networks with full-precision weights in both benchmark tasks. For reference, we consider a spiking network that consolidates weights with activity-dependent metaplasticity (details in Methods) [11]. Since this weight consolidation mechanism requires precise weight updates, we accumulate the weight gradients in auxiliary memory, which is reset after every sample. This model is evaluated with multi-memristor weights ($n_{\text{mem}}$ = 7) considering the same network dimensions. We see that activity-dependent metaplasticity leads to performance similar to probabilistic metaplasticity in benchmark tasks. However, the latter requires $\sim 67\%$ lower memory overhead.
@@
Finally, we perform two controlled experiments to show the contribution of probabilistic metaplasticity for continual learning. First, we study the contribution of consolidating weights with high magnitude and high adjacent neural activity. For this study, we explore random consolidation in which the update probabilities of the weights are calculated with equation 1, but the results are randomly shuffled among the weights. Second, we study the contribution of probabilistic metaplasticity in contrast to probabilistic plasticity. The controlled experiment evaluates this by updating all weights with the same update probability. We also reduce the weight update probability or plasticity by a factor with each incoming task to help retain previous knowledge. Table 2 shows the results of the controlled experiments on the split-MNIST benchmark. We carried out the experiments considering multi-memristor weights ($n_{\text{mem}}$ = 7) in the same spiking network as in the previous experiments. The baseline network is trained based on an error threshold with no additional mechanism. The table lists the accuracies of each task in the benchmark and the mean accuracy across tasks after sequential training. We see that random consolidation fails to learn continually and suffers from catastrophic forgetting. The performance of the decaying probabilistic plasticity varies with the update probability reduction factor. A high reduction factor leads to consolidation in weights, which retains high accuracies in earlier tasks. However, the network fails to learn the last task. A low reduction factor shows the opposite characteristic, where the network catastrophically forgets the earlier tasks. Controlled experiments on the split-Fashion MNIST benchmark show the same response (Supplementary note S2). Contrary to the experiments, we see that probabilistic metaplasticity retains high accuracy in earlier tasks with much lower degradation in the accuracy of the last task compared to decaying probabilistic plasticity. Since the proposed mechanism consolidates weights selectively based on their magnitude and adjacent neural activity, it retains high plasticity in weights assigned low importance, leading to a better balance of stability and plasticity.
@@
Figure 3: Mixed-signal architecture and computational flow for the continual learning mechanisms. We consider a mixed-signal architecture where memristor crossbars carry out the forward pass of the spiking model and the error computation and weight update decisions are carried out by digital circuit modules. The figure shows the computational flow for both the proposed mechanism with probabilistic metaplasticity and the gradient accumulation approach with activity-dependent metaplasticity.
@@
We assess energy consumption for continual learning with probabilistic metaplasticity by analyzing the computations and memory accesses required to train on the benchmark tasks. We then determine the energy to train one sample by analyzing the average number of operations and the individual operation energy requirements (details in Supplementary note S3). We also perform the analysis for activity-dependent metaplasticity with gradient accumulation as a reference. Figure 3 shows the architecture and computational flow for the analysis. We focus on the energy consumption during the parameter update phase since the forward pass and error computations are common across both approaches. We analyze the activities of a spiking network with 200 hidden neurons and multi-memristor weights ($n_{\text{mem}}$ = 7). The mixed-signal and digital components for different network operations were designed in the IBM 65nm 10LPe process. We designed a $16\times 16$ 1T1R crossbar array with peripherals to estimate the energy for reading and writing memristor weights. The metaplastic coefficients and the gradients for the activity-dependent metaplasticity mechanism were modeled as on-chip SRAM with 16-bit and 32-bit bus widths respectively (more details in Supplementary note S3).
Figure 4: Energy consumption during parameter update phase. a. Energy required during the parameter update phase per sample with probabilistic metaplasticity and activity-dependent metaplasticity with gradient accumulation considering exponential (exp) and bilinear (lin) metaplasticity functions. b. Breakdown of energy consumption during parameter update in the split-MNIST task. In the diagram, "Computation" denotes all operations that do not involve SRAM or memristor read/write.

### v5

@@ Abstract
Edge devices operating in dynamic environments critically need the ability to continually learn without catastrophic forgetting. The strict resource constraints in these devices pose a major challenge to achieve this, as continual learning entails memory and computational overhead. Crossbar architectures using memristor devices offer energy efficiency through compute-in-memory and hold promise to address this issue. However, memristors often exhibit low precision and high variability in conductance modulation, rendering them unsuitable for continual learning solutions that require precise modulation of weight magnitude for consolidation. Current approaches fall short to address this challenge directly and rely on auxiliary high-precision memory, leading to frequent memory access, high memory overhead, and energy dissipation. In this research, we propose probabilistic metaplasticity, which consolidates weights by modulating their update probability rather than magnitude. The proposed mechanism eliminates high-precision modification to weight magnitudes and, consequently, the need for auxiliary high-precision memory. We demonstrate the efficacy of the proposed mechanism by integrating probabilistic metaplasticity into a spiking network trained on an error threshold with low-precision memristor weights. Evaluations of continual learning benchmarks show that probabilistic metaplasticity achieves performance equivalent to state-of-the-art continual learning models with high-precision weights while consuming $\sim 67\%$ lower memory for additional parameters and up to 60$\sim$ 60$\times$ 60 × lower energy during parameter updates compared to an auxiliary memory-based solution. The proposed model shows potential for energy-efficient continual learning with low-precision emerging devices.
@@ Introduction
Autonomous edge devices constitute a major application domain of artificial intelligence–ranging from indoor robots to autonomous drones/vehicles [1]. These devices often need on-device training for real-time adaptation to dynamic environments [2]. However, training neural networks requires energy-intensive data movement and memory accesses, which can strain the energy budget of these resource-constrained platforms [3]. The challenge exacerbates in real-world scenarios where data is non-stationary and does not maintain independent and identical distribution. Conventional artificial intelligence fails to learn effectively in this scenario and catastrophically forgets previously learned information [4]. The ability to learn and adapt to non-stationary data streams requires continual learning mechanisms, which entail additional memory and computational overhead, further straining the strict resource budget on edge devices [5]. The constraints in autonomous edge applications present unique challenges, which lead to certain desirable criteria for continual learning solutions. First, the continual learning solution should be task-agnostic, i.e., it should be effective with no supervision or task-boundary awareness for autonomous applications. Second, edge devices have limited on-chip memory, which prohibits numerous iterative training over a large dataset. Hence, these devices should learn in an online scenario where data are encountered only once [1]. Finally, the memory and computational overhead of the continual learning mechanisms should be minimized considering the strict energy constraints of edge devices [6].
@@
Although spiking models incorporated in memristor crossbar architectures are promising, they show non-ideal characteristics which can limit their suitability for continual learning mechanisms. The weight consolidation mechanism mentioned above is a prominent method of continual learning that relies on precise modulation of the weight magnitude. This is extremely challenging to achieve with memristors since these nanoscale devices show low precision ($\sim 3-6$ bits [20, 21]) and inherent variability during conductance modulation. Recent research has circumvented this issue by incorporating auxiliary high-precision memory (usually implemented with Static Random Access Memory (SRAM)) alongside memristor crossbars [22]. Binary memristor devices have been used to accelerate forward pass in continual learning tasks while training with backpropagation is carried out in high-precision memory [23]. Memristors with multi-level conductance have also been utilized for continual learning with metaplasticity, where auxiliary high-precision memory accumulates weight gradients, and memristor weights are updated when the accumulated gradient is equivalent to the device conductance resolution [24]. These approaches have shown promising performance with low-precision memristor weights in continual learning benchmarks. However, these solutions rely on multi-epoch training and knowledge of task progression in order to either assign separate output neurons for different tasks or to modulate hyperparameters. These prerequisites are not suitable for autonomous applications. Moreover, high-precision memory with SRAMs lead to a large silicon footprint and idle standby leakage power (tens to hundreds of picoWatts per bit) [25]. Finally, while crossbar architectures with auxiliary high-precision memory retain energy efficiency during inference, excessive memory access during training can potentially void the benefits of the compute-in-memory architecture.
@@
In this research, we explore a novel weight consolidation mechanism suitable for low-precision memristor weights. Neuroscience literature has shown that binary metaplastic synapses characterized by probabilistic transitions help memory retention [26]. Inspired by this, we propose the probabilistic metaplasticity model, where a weight’s plasticity translates to its probability of update rather than its magnitude of update. The probability of a weight’s update evolves during training such that important weights are assigned lower update probability to prevent catastrophic forgetting. Since this model does not need precise modulation in weight magnitude, the need for auxiliary high-precision memory is eliminated. Previous research explored probabilistic metaplasticity in the context of binary synapses in models of memory [26] and binarized networks that only support training a single-layer network [27]. We generalize this concept to memristor weights with multi-level conductance incorporated in a multilayer spiking network trained based on an error threshold. We evaluate the proposed solution on image detection continual learning benchmarks considering single and multi-memristor weights. Evaluations show that low-precision memristor weights with probabilistic metaplasticity can achieve performance equivalent to state-of-the-art networks with high-precision weights in the challenging online task-agnostic continual learning scenario. Compared to an equivalent solution with high-precision auxiliary memory, probabilistic metaplasticity reduces memory overhead for additional parameters by $\sim 67\%$ while requiring up to 60$\sim$ 60$\times$ 60 × lower energy during the parameter update phase. The proposed model is promising for memory- and energy-efficient continual learning for autonomous edge applications.
@@
Figure 1: Probabilistic metaplasticity with error threshold-based training. a. Spiking network trained on error threshold where the network weights are realized with 1T1R memristor crossbar array. The dendritic compartments in the hidden and output layer neurons integrate the error. When the error reaches the threshold $U_{\text{th}}$, the memristor weights are updated to the next higher (negative error) or lower (positive error) conductance level with probability $p_{\text{update}}$. b. Mean and standard deviation of resistance levels versus programming compliance current for the 1T1R memristor device (shown in inset) adopted in this work [20]. c. Update probability of weights for different values of the metaplasticity coefficient $m$ and the weights $w$. $m$ is positively associated with the activity level of adjacent neurons. We see that weights with highly active adjacent neurons (high $m$) and weight magnitude (high $|w|$) lead to low update probability.
@@
Here, $m_{ij}$ is a metaplasticity coefficient assigned to each weight. As the network learns, the metaplasticity coefficient evolves based on the activity trace of the adjacent neurons as -
$m_{ij}(t+1)=\begin{cases}m_{ij}(t)+{\Delta}m,&\textrm{\leavevmode\nobreak\ if}\>X^{\text{tr}}_{i}\geq m^{\text{pre}}_{\text{th}}\&X^{\text{tr}}_{j}\geq m^{\text{post}}_{\text{th}}\\ m_{ij}(t)&\textrm{\leavevmode\nobreak\ otherwise}\end{cases}.$ (2)
We see that high activity in the adjacent neurons increases the metaplasticity coefficient of a weight. Figure 1c shows the update probability for different values of weights and metaplasticity coefficients. We see that weights with higher magnitude and metaplasticity coefficient exhibit lower update probability. Throughout training, a weight’s magnitude and metaplasticity coefficient adapt in response to the activity of the network. This changes the update probability/plasticity of a weight, leading to a metaplastic effect to consolidate weights.
@@
                                     if $\epsilon<p_{\text{update}}$ then
@@
            Update metaplasticity coefficient: $m\leftarrow m+{\Delta}m$
@@
Algorithm 1 Probabilistic metaplasticity with error threshold-based eRBP for continual learning in a spiking network. $\mathcal{T}$ denotes a set of sequential tasks. $S^{\text{in}}$ and $S^{\text{out}}$ are the input and output spike trains and $W$ denotes weights realized with memristors. $U$ is the error accumulated at the dendritic compartment of neurons. When |$U$| crosses the error threshold $U_{\text{th}}$, the update probability $p_{\text{update}}$ of eligible weights are calculated and compared with a random number to decide which weights should be updated. The memristor weights selected for update are programmed to the next higher or lower conductance level depending on the sign of the error. The function Program() refers to the operations required to update a memristor’s conductance. Details of the eRBP algorithm can be found in methods.
@@ Results
We first evaluate probabilistic metaplasticity on the split-MNIST benchmark, which consists of five sequential image classification tasks, each composed of two classes from the MNIST dataset [34]. We conduct network-level simulations in Python considering a spiking network with 200 hidden neurons and 2 output neurons. The simulation emulates in situ learning with memristor weights and captures the quantized nonuniform distribution and the associated variance of memristor conductance during weight updates. Figure 2 shows the evolution of the network performance in each of the benchmark tasks as the network trains on them sequentially. When trained with error threshold-based eRBP with no probabilistic metaplasticity, we see that the network learns each task it encounters with high accuracy (Figure 2a). However, after training on all tasks, it retains high accuracy only on the recently encountered tasks, and the earlier task accuracies drop to random guess level.
Figure 2: Evolution of task accuracies with sequential training on the split-MNIST benchmark. The x-axis shows the latest task the network has learned. We evaluate the performance of task $n$ only after the network encounters it and denote the accuracies as 0 before that. a. With no probabilistic metaplasticity, the network learns the current task well, but forgets the initial tasks after sequentially learning multiple tasks. b. Probabilistic metaplasticity with low activity threshold leads to high rigidity in the network, so it remembers previous tasks but cannot learn the last task. c. High activity threshold can lead to loss in previous task accuracy while the network remains plastic to learn the new tasks. d. Optimized activity threshold balances plasticity and rigidity such that the network maintains high initial task accuracies while maintaining the ability to learn new tasks.
@@
We further evaluate probabilistic metaplasticity on the split-Fashion MNIST benchmark, which also consists of five sequential image classification tasks, each with two classes [35]. We compare the performance of probabilistic metaplasticity with state-of-the-art models which also consolidate important weights or regularize weight updates to achieve continual learning. These approaches differ in the criteria that determine weight importance. “Synaptic Intelligence” (SI) [13] measures weight importance based on the contribution of weights to improve network performance, whereas “Memory Aware Synapses” (MAS) [36] measures it through the sensitivity of the network’s output function to changes in weights. “Stochastic Synapses” (SS) [37] adopts weights which transmit information probabilistically and weights with high transmission probability are consolidated. “Bayesian Gradient Descent” (BGD) [38] is a Bayesian continual learning mechanism that assesses the importance of parameters based on the uncertainty associated with their distribution. “Learning without forgetting” (LwF) [39] preserves the response of the network to previous tasks through knowledge distillation. “TACOS” [11] introduces activity-dependent metaplasticity where weights are consolidated based on their magnitude and the activity level of adjacent neurons. Except for TACOS, which was evaluated in a spiking network, the rest of the models were evaluated considering multi-layer perceptrons. Weight consolidation models typically require additional memory to store information for consolidating weights, which can be implemented with on-chip memory in hardware. In this work, we model this memory as on-chip SRAM during the energy analysis in the following subsection. Table 1 shows the memory required by each continual learning model for parameters, excluding weights, considering the same network architecture used to evaluate probabilistic metaplasticity (see Supplementary note S1 for details). We exclude the network weights in the memory overhead computation since the aforementioned continual learning models were evaluated considering full-precision weights, whereas the proposed mechanism realizes weights with memristor devices. We also evaluate a spiking neural network with full-precision weights trained with eRBP as a baseline.
@@
Table 1 lists the mean accuracy across tasks for the different models after sequentially training on multiple tasks. To observe the effect of weight resolutions on the performance of probabilistic metaplasticity, we consider both single- and multi-memristor weights[40]. Multi-memristor weights are configured by connecting $n_{\text{mem}}$ number of memristors in parallel. Figure 1a shows a 1T1R crossbar with multi-memristor weights consisting of three 1T1R devices. During training, only one memristor arbitrated by a global counter is updated to reduce training overhead. The parallel combination of devices increases the dynamic range of the weight conductance by a factor of $n_{\text{mem}}$ and increases the weight resolution. We note an improvement in the mean accuracy across tasks with increased weight resolution, with the best performance when $n_{\text{mem}}$ = 7, leading to a mean weight resolution $\sim 6$-bit. Results show that with probabilistic metaplasticity, a spiking network with noisy low-precision multi-memristor weights achieves performance equivalent to state-of-the-art networks with full-precision weights in both benchmark tasks. For reference, we consider a spiking network that consolidates weights with activity-dependent metaplasticity (details in Methods) [11]. Since this weight consolidation mechanism requires precise weight updates, we accumulate the weight gradients in auxiliary memory, which is reset after every sample. This model is evaluated with multi-memristor weights ($n_{\text{mem}}$ = 7) considering the same network architecture as probabilistic metaplasticity. We see that activity-dependent metaplasticity leads to performance similar to probabilistic metaplasticity in benchmark tasks. However, the latter requires $\sim 67\%$ lower memory overhead. We extend the performance analysis by evaluating probabilistic metaplasticity on the split-CIFAR-10 benchmark [41], which consists of five sequential image detection tasks. It is a more complex dataset compared to the previous benchmarks, containing larger images, more complex image types and colored images instead of grayscale. To address this complexity, we consider a ResNet-18 backbone pre-trained on ImageNet to process the images and use the extracted features as inputs to the spiking networks [42, 32]. Probabilistic metaplasticity was evaluated considering a spiking network with 200 hidden neurons and multi-memristor weights ($n_{mem}$ =7). We also evaluate a baseline spiking network with multi-memristor weights and a spiking network trained with TACOS. Table 3 lists the accuracies of individual tasks and mean accuracy across tasks. We see that although probabilistic metaplasticity adopts low-precision multi-memristor weights, it demonstrates performance equivalent to TACOS, which adopts full-precision weights.
We further analyze the robustness of the proposed probabilistic metaplasticity model by evaluating its performance on a smaller set of training data. An autonomous agent learning in an online setting may need to learn from fewer samples. To explore this scenario, we trained and optimized the model on a subset of the split-MNIST dataset and tested it on the complete test set. We observed some degradation in the mean accuracy across tasks with reduced dataset size (shown in Figure 3. The degradation may stem from the probabilistic consolidation mechanism’s reliance on the law of large numbers. However, we observed a similar degradation when training the network with activity-dependent metaplasticity. This suggests that the decline in performance could also be attributed to the general performance degradation in machine learning models when the number of training samples is insufficient.
Figure 3: Effect of number of training samples on continual learning. The mean accuracy across tasks shows degradation as the average number of samples per class reduces.
@@
Finally, we perform two controlled experiments to show the contribution of probabilistic metaplasticity for continual learning. First, we study the contribution of consolidating weights with high magnitude and high adjacent neural activity. For this study, we explore random consolidation in which the update probabilities of the weights are calculated with equation 1, but the results are randomly shuffled among the weights. Second, we study the contribution of probabilistic metaplasticity in contrast to probabilistic plasticity. The controlled experiment evaluates this by updating all weights with the same update probability. We also reduce the weight update probability or plasticity by a factor with each incoming task to help retain previous knowledge. Table 2 shows the results of the controlled experiments on the split-MNIST benchmark. We carried out the experiments considering multi-memristor weights ($n_{\text{mem}}$ = 7) in the same spiking network as in the previous experiments. The baseline network is trained based on an error threshold with no additional mechanism. The table lists the accuracies of each task in the benchmark and the mean accuracy across tasks after sequential training. We see that random consolidation fails to learn continually and suffers from catastrophic forgetting. The performance of the decaying probabilistic plasticity varies with the update probability reduction factor. A high reduction factor leads to consolidation in weights, which retains high accuracies in earlier tasks. However, the network fails to learn the last task. A low reduction factor shows the opposite characteristic, where the network catastrophically forgets the earlier tasks. Controlled experiments on the split-Fashion MNIST benchmark show the same response (Supplementary note S2 and Supplementary Table S2). Contrary to the experiments, we see that probabilistic metaplasticity retains high accuracy in earlier tasks with much lower degradation in the accuracy of the last task compared to decaying probabilistic plasticity. Since the proposed mechanism consolidates weights selectively based on their magnitude and adjacent neural activity, it retains high plasticity in weights assigned low importance, leading to a better balance of stability and plasticity.
@@
Figure 4: Mixed-signal architecture and computational flow for the continual learning mechanisms. We consider a mixed-signal architecture where memristor crossbars carry out the forward pass of the spiking model and the error computation, arithmetic and logical operations are carried out in the digital domain. The figure shows the computational flow for continual learning with probabilistic metaplasticity (left) and activity-dependent metaplasticity with gradient accumulation (right).
@@
We evaluate energy consumption for continual learning with probabilistic metaplasticity by analyzing the computations and memory accesses required to train on the benchmark tasks. We then determine the energy needed to train on a sample by analyzing the average number of operations and the individual operation energy requirements (details in Supplementary note S3). We also perform the analysis for activity-dependent metaplasticity with gradient accumulation as a reference. Figure 4 shows the architecture and computational flow for the analysis. We focus on the energy consumption during the parameter update phase (the operations in the gray box in Figure 4) since the operations incurred during the forward pass and error computations are the same in both approaches. We analyze the activities of a spiking network with 200 hidden neurons and multi-memristor weights ($n_{\text{mem}}$ = 7). The mixed-signal and digital components for different network operations were designed in the IBM 65nm 10LPe process. We designed a $16\times 16$ 1T1R crossbar array with peripherals to estimate the energy to read and write memristor weights. The metaplasticity coefficients for both mechanisms and the gradients for the activity-dependent metaplasticity mechanism were modeled as on-chip SRAM with 16-bit and 32-bit bus widths, respectively (more details in Supplementary note S3).